OpenStudy (anonymous):

MEDAL REWARDED Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x.

OpenStudy (anonymous):

f(x)=

OpenStudy (anonymous):

g(x)=

hero (hero):

I can help with these because these are my favorite things to work on

hero (hero):

What I usually do first is put one of the functions in mixed form.

OpenStudy (anonymous):

I understand the concept of what im supposed to do, substitute the x and make sure its equals the same thing on both sides, but i get mixed up and end up confusing myself.

hero (hero):

Yeah, I know, but if you just plug it in like that, it will make it more difficult to reduce.

OpenStudy (anonymous):

Okay, show me how you do it your way then :)

hero (hero):

\[f(x) \\= \frac{x - 9}{x + 5} \\= \frac{x + 5 - 14}{x + 5} \\= \frac{x + 5}{x + 5} - \frac{14}{x + 5} \\= 1 - \frac{14}{x + 5}\]

hero (hero):

Oh look, now we only have one x

hero (hero):

Now we can insert g(x) in to the one x rather than insert g(x) in to two x's

OpenStudy (anonymous):

how does it look once you add g(x) in?

OpenStudy (anonymous):

using your way.

hero (hero):

But let's see if we can make g(x) easier to deal with: \[g(x) = \frac{-5x - 9}{x - 1} = \frac{-(5x + 9)}{x - 1}\]

hero (hero):

I had it completely typed out and I messed up and accidentally erased everythign I typed

hero (hero):

Now you can insert g(x) into f(x) and get what you need

hero (hero):

Let me demonstrate for you

OpenStudy (anonymous):

Im sorry but im confused now. I would appreciate it if you could do everything at once? If not its fine. I know its alot of work

OpenStudy (anonymous):

@Hero ?

hero (hero):

Sorry, I had to re-start my computer. I will show you

OpenStudy (anonymous):

Okay thank you, this is my last problem on my review

OpenStudy (anonymous):

Are you there? it wont let me tag you

hero (hero):

I'm here. I think I made a mistake with one of these. Hang on

OpenStudy (anonymous):

Okay, take your time

hero (hero):

Okay, I see my mistake \[g(x) = -5 - \frac{14}{x-1}\]

hero (hero):

And now we can do f(g(x)) and g(f(x)) relatively easy

hero (hero):

\[f(g(x)) = 1 - \frac{14}{-\frac{14}{x-1} - 5 + 5} \\= 1 - \frac{14}{-\frac{14}{x-1}} \\= 1 + \frac{14}{\frac{14}{x - 1}} \\= 1 + \left(14 \div \frac{14}{x - 1}\right) \\= 1 + \left(14 \times \frac{x - 1}{14}\right) \\=1 + (x - 1) \\= 1 + x - 1 \\=1 - 1 + x \\ = x\]

hero (hero):

I know it doesn't appear that way, but this is the easy way.

hero (hero):

You can do the same thing with \(g(f(x))\)

OpenStudy (anonymous):

I only need to prove that one side equals x because then the other side obviously does as well. Thank you so much, im going to look over the way you do it now

hero (hero):

I should have color coded the g(x) I inserted into f(x)

hero (hero):

Basically \(f(x) = 1 - \dfrac{14}{x + 5}\) \(g(x) = -5 - \dfrac{14}{x - 1}\)

hero (hero):

When you have fractional functions of this type you only want to insert one into the other ONCE.

hero (hero):

If you have any questions, NOW is the time to ask.

hero (hero):

I suppose you understand it now?